OpenStudy (anonymous):

How do you find the common difference in an arithmetic sequence? Here's an example: Identify the 31st term of an arithmetic sequence where a1 = 26 and a22 = -226.

hartnn (hartnn):

use the n'th term formula \(\Large a_n = a_1 +(n-1)d\) here, n = 22 a1 = 26 and a22 = -226 just plug them in! only 'd' = common difference is unknown

OpenStudy (anonymous):

yeah, I know how to do that.. it's just the common difference that throws me off. how do you find that?

hartnn (hartnn):

\(-226 = 26 + (22-1) d\) find 'd' from here thats your common difference :)

OpenStudy (anonymous):

wait I thought the equation would be: n=31 a1= 26

OpenStudy (anonymous):

so it would a31=26+(31-1)d

hartnn (hartnn):

thats correct equation, but you will need to find 'd' first for that , isn't it ?

OpenStudy (anonymous):

yeah, that's what was throwing me off

hartnn (hartnn):

\(-226 = 26 + (22-1) d\) please first find 'd' from here...

hartnn (hartnn):

i got that equation using the n'th term formula only

OpenStudy (anonymous):

I got -4.8

hartnn (hartnn):

is it ? 21d= -226 -26 i get d =12.....

OpenStudy (anonymous):

oh I didn't switch around the equation. I got that now

hartnn (hartnn):

now that you have d=12 use the equation you have for a31 :)

hartnn (hartnn):

oh, and sorry, d should be -12 \(\Large d = -12 \)

OpenStudy (anonymous):

let me work it out see if I have anymore questions

OpenStudy (anonymous):

i got -334, which is one of my answer choices. thank you for helping me figure this out!!

hartnn (hartnn):

-334 is correct :) and welcome ^_^

hartnn (hartnn):

while you're learning this topic, you can go through this....might be helpful :) http://openstudy.com/study#/updates/503bb2a0e4b007f9003103b0